# Comprehensive Assessment of Jailbreak Attacks Against LLMs

### v2

@@ Introduction
Method Vicuna ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 PaLM2 Average
AIM 0.99(±0.01) 0.93(±0.01) 0.13(±0.05) 0.00(±0.01) 0.00(±0.00) 0.99(±0.01) 0.62(±0.04) 0.88(±0.02) 0.57(±0.42)
Devmoderanti 0.91(±0.02) 0.79(±0.04) 0.14(±0.01) 0.02(±0.01) 0.00(±0.00) 0.73(±0.03) 0.08(±0.03) 0.61(±0.03) 0.41(±0.36)
Devmode v2 0.89(±0.04) 0.65(±0.06) 0.20(±0.01) 0.00(±0.01) 0.00(±0.00) 0.53(±0.04) 0.51(±0.05) 0.54(±0.02) 0.41(±0.30)
Base64 0.15(±0.02) 0.02(±0.00) 0.11(±0.01) 0.00(±0.01) 0.01(±0.02) 0.14(±0.03) 0.49(±0.05) 0.01(±0.01) 0.12(±0.15)
Combination 0.13(±0.01) 0.09(±0.01) 0.06(±0.01) 0.15(±0.02) 0.21(±0.07) 0.31(±0.04) 0.74(±0.04) 0.04(±0.01) 0.22(±0.22)
Zulu 0.18(±0.02) 0.04(±0.01) 0.08(±0.01) 0.14(±0.04) 0.21(±0.01) 0.79(±0.03) 0.76(±0.06) 0.01(±0.01) 0.28(±0.30)
DrAttack 0.85(±0.08) 0.63(±0.03) 0.45(±0.08) 0.35(±0.02) 0.32(±0.01) 0.80(±0.03) 0.79(±0.11) 0.73(±0.08) 0.62(±0.20)
AutoDAN 0.98(±0.01) 0.90(±0.03) 0.58(±0.04) 0.52(±0.07) 0.50(±0.08) / / / 0.70(±0.20)
GPTFuzz 0.79(±0.04) 0.88(±0.05) 0.41(±0.02) 0.31(±0.09) 0.25(±0.04) 0.85(±0.01) 0.41(±0.02) 0.48(±0.01) 0.55(±0.24)
LAA 1.00(±0.01) 0.93(±0.11) 0.88(±0.04) 0.88(±0.06) 0.55(±0.05) 1.00(±0.05) 0.74(±0.06) 0.85(±0.09) 0.85(±0.14)
GCG 0.87(±0.05) 0.44(±0.07) 0.56(±0.04) 0.51(±0.03) 0.48(±0.09) / / / 0.57(±0.15)
COLD 0.50(±0.02) 0.50(±0.09) 0.45(±0.05) 0.41(±0.03) 0.40(±0.09) / / / 0.45(±0.04)
PAIR 0.76(±0.05) 0.54(±0.07) 0.48(±0.04) 0.46(±0.02) 0.41(±0.03) 0.62(±0.05) 0.80(±0.03) 0.78(±0.02) 0.61(±0.15)
TAP 0.74(±0.05) 0.76(±0.04) 0.44(±0.03) 0.47(±0.02) 0.43(±0.07) 0.81(±0.03) 0.71(±0.03) 0.74(±0.02) 0.64(±0.15)
MasterKey 0.88(±0.04) 0.82(±0.05) 0.11(±0.03) 0.07(±0.05) 0.05(±0.03) 0.90(±0.02) 0.54(±0.03) 0.76(±0.03) 0.52(±0.35)
AdvPrompter 0.58(±0.08) 0.50(±0.08) 0.32(±0.08) 0.15(±0.05) 0.17(±0.01) / / / 0.34(±0.17)
Generation Exploitation 0.95(±0.07) 0.80(±0.06) 0.72(±0.08) 0.50(±0.04) 0.44(±0.04) / / / 0.68(±0.19)
Average 0.72(±0.29) 0.60(±0.30) 0.36(±0.24) 0.29(±0.24) 0.26(±0.19) 0.71(±0.25) 0.60(±0.20) 0.54(±0.32) /
Baseline 0.52(±0.03) 0.38(±0.00) 0.31(±0.07) 0.39(±0.13) 0.39(±0.04) 0.44(±0.07) 0.38(±0.08) 0.47(±0.06) 0.41(±0.06)
@@
Table 3: Average ASRs of all jailbreak attacks (direct attack). The baseline here refers to the average ASRs across different violation categories on the eight LLMs without jailbreak attacks. The highest value in a row is highlighted in blue, and the highest value in a column is bolded.
Violation Category Vicuna ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 PaLM2 Average Baseline
Hate, Unfairness or Harassment 0.63(±0.36) 0.52(±0.32) 0.14(±0.16) 0.12(±0.18) 0.09(±0.11) 0.62(±0.31) 0.44(±0.30) 0.46(±0.30) 0.38(±0.21) 0.06(±0.07)
Malicious Software 0.69(±0.30) 0.55(±0.33) 0.30(±0.26) 0.21(±0.24) 0.19(±0.17) 0.65(±0.30) 0.38(±0.29) 0.55(±0.36) 0.44(±0.18) 0.15(±0.15)
Well-being Infringement 0.78(±0.33) 0.71(±0.34) 0.57(±0.36) 0.41(±0.37) 0.41(±0.36) 0.78(±0.36) 0.84(±0.25) 0.64(±0.38) 0.64(±0.16) 0.79(±0.14)
Physical Harm 0.68(±0.38) 0.54(±0.34) 0.19(±0.20) 0.22(±0.24) 0.19(±0.14) 0.61(±0.29) 0.38(±0.23) 0.38(±0.26) 0.40(±0.18) 0.10(±0.12)
Disinformation Spread 0.71(±0.33) 0.64(±0.35) 0.27(±0.27) 0.21(±0.23) 0.15(±0.16) 0.72(±0.27) 0.53(±0.30) 0.54(±0.34) 0.47(±0.21) 0.08(±0.11)
Privacy Breach 0.69(±0.32) 0.52(±0.32) 0.21(±0.23) 0.21(±0.22) 0.16(±0.15) 0.66(±0.28) 0.34(±0.21) 0.51(±0.33) 0.41(±0.20) 0.08(±0.11)
Adult Content 0.78(±0.33) 0.70(±0.33) 0.48(±0.32) 0.37(±0.35) 0.38(±0.30) 0.83(±0.27) 0.83(±0.26) 0.51(±0.33) 0.61(±0.18) 0.83(±0.17)
Political Activities 0.78(±0.32) 0.77(±0.33) 0.58(±0.35) 0.43(±0.37) 0.44(±0.35) 0.85(±0.23) 0.89(±0.19) 0.61(±0.36) 0.67(±0.17) 0.86(±0.16)
Impersonation 0.71(±0.35) 0.68(±0.33) 0.49(±0.36) 0.43(±0.36) 0.36(±0.37) 0.73(±0.30) 0.85(±0.22) 0.60(±0.34) 0.61(±0.16) 0.88(±0.15)
Terrorist Content 0.68(±0.31) 0.40(±0.32) 0.16(±0.22) 0.12(±0.14) 0.09(±0.11) 0.58(±0.36) 0.24(±0.30) 0.48(±0.35) 0.34(±0.21) 0.08(±0.10)
Unauthorized Practice 0.77(±0.33) 0.66(±0.31) 0.64(±0.31) 0.47(±0.34) 0.40(±0.36) 0.76(±0.33) 0.83(±0.24) 0.58(±0.39) 0.64(±0.14) 0.78(±0.13)
Safety Filter Bypass 0.72(±0.28) 0.56(±0.32) 0.39(±0.29) 0.33(±0.29) 0.31(±0.29) 0.69(±0.24) 0.66(±0.28) 0.53(±0.34) 0.52(±0.15) 0.26(±0.19)
Risky Government Decisions 0.69(±0.34) 0.61(±0.34) 0.18(±0.21) 0.18(±0.17) 0.19(±0.15) 0.67(±0.30) 0.45(±0.24) 0.63(±0.38) 0.45(±0.22) 0.34(±0.27)
AI Usage Disclosure 0.79(±0.33) 0.75(±0.33) 0.56(±0.41) 0.42(±0.38) 0.41(±0.37) 0.81(±0.23) 0.85(±0.20) 0.55(±0.36) 0.64(±0.17) 0.94(±0.07)
Third-party Rights Violation 0.70(±0.30) 0.55(±0.31) 0.41(±0.28) 0.29(±0.29) 0.25(±0.26) 0.75(±0.28) 0.67(±0.21) 0.54(±0.34) 0.52(±0.18) 0.29(±0.15)
Illegal Activities 0.62(±0.34) 0.46(±0.37) 0.22(±0.23) 0.19(±0.16) 0.16(±0.15) 0.62(±0.33) 0.43(±0.29) 0.46(±0.27) 0.39(±0.17) 0.03(±0.04)
@@
Table 2 shows the ASR results obtained from all the target LLMs and jailbreak attacks based on our taxonomy. We observe that none of the eight LLMs demonstrate initial complete resistance to harmful questions. For the well-aligned LLMs such as Llama2, Llama3, and GPT-4, the baseline ASRs are 0.31, 0.39, and 0.38, respectively. The average ASR of most LLMs exceeds the baseline, except Llama3 and Llama3.1, indicating the overall vulnerability of most LLMs to jailbreak attacks. Among all the models, Vicuna is the most vulnerable to jailbreak attacks, which aligns with our conclusion related to the unified policy (see Section 6.2) and previous works [81, 66, 49].
From our analysis of the jailbreak taxonomy, we observe that jailbreak methods classified under obfuscation-based approaches generally yield the lowest attack performance, as outlined in Table 2. Notably, DrAttack emerges as an exception. After detailed analysis, we discern that most obfuscation-based jailbreak attacks are model-specific attacks, i.e., they can achieve a high ASR score on a specific model, especially on GPT-4. For example, Zulu could only achieve good attack performance on GPT-3.5 and GPT-4, with the ASR exceeding 0.75. We attribute this phenomenon to the strong capability of the LLM. LLMs like GPT-4, with their extensive training on diverse datasets, obtain the capacity to recognize and process texts that have been translated to low-resource languages or encoded, which is beyond the capability of other models. However, this capability is a double-edged sword, as it also enlarges the attack surface of the high-performance LLMs. The extra attack surface makes it harder to align such LLMs and increases their vulnerability to jailbreak attacks. Thus, the vulnerability is exploited by obfuscation-based jailbreak methods, and these methods could achieve significant attack performance against the specific LLMs. On the other hand, the semantic-based vulnerability exploited by DrAttack is cross-model, showcasing the broad applicability and potential of this approach.
In general, from Table 2, AutoDAN and Generation Exploitation achieve the highest ASR (above 0.68) when targeting all LLMs in the white-box setting. AdvPrompter performs worse in the white-box setting, which may be because we use a relatively small LLM (TinyLlama-1.1B-step-50K-105b), and the distribution of the test set is significantly different from its fine-tuning dataset.
Among the black-box scenarios, LAA achieves the best attack performance. To illustrate, LAA achieves 0.85 average ASR scores—the highest among all the jailbreak methods. It even achieves ASRs exceeding 0.55 on all tested LLMs. Notably, for the black-box scenario, we find that the human-based jailbreak prompts (also referred to as “in the wild” prompts) can still achieve high ASR in many cases, which highlights the significance of actively collecting and analyzing such jailbreak prompts.
@@
The results in Table 3 show that the ASR scores vary significantly across different violation categories from our unified policy. In general, Hate, Unfairness or Harassment, Physical Harm, Terrorist Content, and Illegal Activities are the three most challenging categories to launch jailbreak attacks, whereas the Political Activities, Well-being Infringement, and AI Usage Disclosure are more straightforward and amenable to jailbreak attacks. To better understand the results of our policy, combined with Table 9, we conduct a fine-grained evaluation, analyzing the performance of various LLMs with our unified policy. Our results reveal that despite the usage policy from the model provider explicitly stating the coverage of violation categories, the ASR results continue to register a high score. For example, despite OpenAI’s explicit prohibition of Political Activities in their policy, the jailbreak attack within this category still manages to achieve the highest ASR (over 0.80) on GPT-3.5 and GPT-4. Meta and Google also experience similar situations. This phenomenon underscores the challenges associated with effectively aligning LLM policies and their jailbreak attacks, indicating the future works for improving LLM policy compliance. In addition, some models, such as Vicuna and ChatGLM3, originally did not provide any detailed usage policy. To mitigate the effectiveness of jailbreak attacks, we expect these model trainers to carry out targeted optimizations in subsequent updates according to our policy-based experimental statistics.
We also identify the LLM with the highest ASR score for each violation category under different jailbreak attacks. Our experimental results indicate that only three LLMs contain the highest ASR scores corresponding to certain violation categories (blue texts in Table 3): Vicuna (8 categories), GPT-3.5 (4 categories), and GPT-4 (4 categories). Therefore, we believe that although other models also suffer from jailbreak attacks across different violation categories, these three models are the most vulnerable. The potential reason is that for most attacks [76, 77, 72], their original targets are the LLMs of OpenAI, i.e., GPT-3.5 and GPT-4. As for the Vicuna, it does not implant any specific safeguards during the fine-tuning and the dataset used is collected from ShareGPT [^4] https://sharegpt.com/. and has not been rigorously reviewed.

### v3

@@ Introduction
Method Vicuna ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 DeepSeek-V3 PaLM2 Average
AIM 0.99 0.93 0.13 0.00 0.00 0.99 0.62 1.00 0.88 0.62
Devmoderanti 0.91 0.79 0.14 0.02 0.00 0.73 0.08 0.56 0.61 0.43
Devmode v2 0.89 0.65 0.20 0.00 0.00 0.53 0.51 0.52 0.54 0.43
Base64 0.15 0.02 0.11 0.00 0.01 0.14 0.49 0.49 0.01 0.16
Combination 0.13 0.09 0.06 0.15 0.21 0.31 0.74 0.78 0.04 0.28
Zulu 0.18 0.04 0.08 0.14 0.21 0.79 0.76 0.49 0.01 0.30
DrAttack 0.85 0.63 0.45 0.35 0.32 0.80 0.79 0.74 0.73 0.63
AutoDAN 0.98 0.90 0.58 0.52 0.50 / / / / 0.70
GPTFuzz 0.79 0.88 0.41 0.31 0.25 0.85 0.41 0.79 0.48 0.58
LAA 1.00 0.93 0.88 0.88 0.55 1.00 0.74 1.00 0.85 0.87
GCG 0.87 0.44 0.56 0.51 0.48 / / / / 0.57
COLD 0.50 0.50 0.45 0.41 0.40 / / / / 0.45
PAIR 0.76 0.54 0.48 0.46 0.41 0.62 0.80 0.92 0.78 0.64
TAP 0.74 0.76 0.44 0.47 0.43 0.81 0.71 0.76 0.74 0.65
Masterkey 0.88 0.82 0.11 0.07 0.05 0.90 0.54 0.95 0.76 0.56
AdvPrompter 0.58 0.50 0.32 0.15 0.17 / / / / 0.34
GE 0.95 0.80 0.72 0.50 0.44 / / / / 0.68
Average 0.72 0.60 0.36 0.29 0.26 0.71 0.60 0.75 0.54 /
Baseline 0.52 0.38 0.31 0.39 0.39 0.44 0.38 0.49 0.47 0.42
@@
Table 3: Average ASRs of all jailbreak attacks (direct attack) across different violation categories. The highest value in a row (not including baseline) is in blue, and the highest value in a column is bolded.
Violation Category Vicuna ChatGLM3 Llama2 Llama3 Llama3.1 GPT-3.5 GPT-4 DeepSeek-V3 PaLM2 Average Baseline Illegal Activities 0.62 0.46 0.22 0.19 0.16 0.62 0.43 0.58 0.46 0.42 0.03 Hate, Unfairness or Harassment 0.63 0.52 0.14 0.12 0.09 0.62 0.44 0.56 0.46 0.40 0.06 Terrorist Content 0.68 0.40 0.16 0.12 0.09 0.58 0.24 0.56 0.48 0.37 0.08 Disinformation Spread 0.71 0.64 0.27 0.21 0.15 0.72 0.53 0.67 0.54 0.49 0.08 Privacy Breach 0.69 0.52 0.21 0.21 0.16 0.66 0.34 0.54 0.51 0.43 0.08 Physical Harm 0.68 0.54 0.19 0.22 0.19 0.61 0.38 0.62 0.38 0.42 0.10 Malicious Software 0.69 0.55 0.30 0.21 0.19 0.65 0.38 0.70 0.55 0.47 0.15 Safety Filter Bypass 0.72 0.56 0.39 0.33 0.31 0.69 0.66 0.83 0.53 0.56 0.26 Third-party Rights Violation 0.70 0.55 0.41 0.29 0.25 0.75 0.67 0.68 0.54 0.54 0.29 Risky Government Decisions 0.69 0.61 0.18 0.18 0.19 0.67 0.45 0.74 0.63 0.48 0.34 Unauthorized Practice 0.77 0.66 0.64 0.47 0.40 0.76 0.83 0.88 0.58 0.67 0.78 Well-being Infringement 0.78 0.71 0.57 0.41 0.41 0.78 0.84 0.95 0.64 0.67 0.79 Adult Content 0.78 0.70 0.48 0.37 0.38 0.83 0.83 0.90 0.51 0.64 0.83 Political Activities 0.78 0.77 0.58 0.43 0.44 0.85 0.89 0.98 0.61 0.70 0.86 Impersonation 0.71 0.68 0.49 0.43 0.36 0.73 0.85 0.92 0.60 0.64 0.88 AI Usage Disclosure 0.79 0.75 0.56 0.42 0.41 0.81 0.85 0.88 0.55 0.67 0.94
  • ” denotes that the violation category is consistently labeled as “general” violations by three human annotators.
@@
Table 2 presents ASR results of different jailbreak attacks. We observe that none of the eight LLMs demonstrate initial complete resistance to forbidden questions. Even for the well-aligned LLMs such as Llama3, the baseline ASR is 0.39.101010We discuss results of the baseline in Appendix D.2. All LLMs suffer from jailbreak attacks, with ASRs exceeding 0.55 under at least one attack method. Notably, the latest model we test, DeepSeek-V3, suffers from the highest average ASR value (0.75), indicating that the jailbreak risk does not have high priority for some developers.
Human-based methods perform well in most cases; however, on certain strongly safety-aligned models (e.g., the Llama3 series), ASR degrades to nearly zero. This is likely because these highly aligned models internally implement rules to detect and reject such static and non-diverse human-based jailbreak attacks. Many other jailbreak methods, such as MasterKey and GPTFuzz, use human-based methods as their initial seeds. As a result, their outcomes exhibit similar trends.
Most obfuscation-based attacks, except DrAttack, are model-specific, often performing well on high-capability models like GPT-4 and DeepSeek-V3. For instance, Zulu achieves ASRs exceeding 0.75 only on GPT-3.5 and GPT-4. This may stem from the advanced abilities of models like GPT-4, trained on diverse datasets, to process low-resource languages or encoded texts—capabilities lacking in other models. However, this also expands their attack surface, making alignment harder and increasing vulnerability to jailbreaks. DrAttack exploits cross-model semantic vulnerabilities, demonstrating broader applicability.
Feedback-based methods do not exhibit significant weaknesses and perform relatively stably, with no extreme cases where the ASR falls below 0.40.
GE, despite querying only with the original forbidden question, achieves a considerable average ASR of 0.68, ranking third among all methods.
LAA outperforms all other attacks, including those white-box attacks, achieving 0.87 average ASR. It even obtains an ASR reaching 0.55 on the safest Llama3.1. This result underscores the reality and urgency of jailbreak risks: even in the most realistic black-box scenarios, highly effective jailbreak attacks exist, making it highly possible for LLMs to be misused.
@@
The results in Table 3 show significant variation in ASRs across violation categories under our unified policy. We observe six specific violation categories (Well-being Infringement & Adult Content & Political Activities & Impersonation & Unauthorized Practice & AI Usage Disclosure), even some of them being explicitly covered in the providers’ usage policies, have higher ASRs on both baseline and average values of all jailbreak attacks than other categories. For instance, OpenAI explicitly prohibits Political Activities, yet this category achieves the highest ASR ($\geq$ 0.80) on GPT-3.5 and GPT-4, with similar results for Meta and Google.
Categories labeled as “general” (Hate, Unfairness or Harassment, Physical Harm, Terrorist Content, and Illegal Activities) are all challenging for jailbreaking, showing the effort of model providers to align LLMs with human preference. Although Disinformation Spread and Privacy Breach are not consistently labeled as “general,” they are still relatively difficult to jailbreak, with average ASR values of only 0.49 and 0.43, respectively.
We identify the LLMs with the highest ASRs for each violation category under different jailbreak attacks. The results show that only three models contain the highest ASR scores across categories (blue texts in Table 3): Vicuna (5 categories), GPT-3.5 (3 categories), and DeepSeek-V3 (9 categories). This indicates that, while other models are also suffering from jailbreaking, these three are the most susceptible. One possible reason is that most attacks [88, 89, 82] are originally designed to target OpenAI’s models. And the three most vulnerable LLMs are either from OpenAI or trained on ChatGPT’s output [20, 92].